# Deep Neural Networks and Brain Alignment: Brain Encoding and Decoding (Survey)

### v2

@@ 2 Stimulus Representations
In this section, we discuss types of stimulus representations proposed in the literature across different modalities: text, visual, audio, video, and other multimodal stimuli.
Text stimulus representations. Older methods for text-based stimuli representation include text corpus co-occurrence counts (Mitchell et al., 2008; Pereira et al., 2013; Huth et al., 2016), topic models (Pereira et al., 2013), syntactic features and discourse features (Wehbe et al., 2014). In recent times, for text-based stimuli, both semantic models and experiential attribute models have been explored. Semantic representation models include word embedding methods (Pereira et al., 2018; Wang et al., 2020; Pereira et al., 2016; Toneva & Wehbe, 2019; Anderson et al., 2017a; Oota et al., 2018), sentence representation models (Sun et al., 2020; 2019; Toneva & Wehbe, 2019), RNNs (Jain & Huth, 2018; Oota et al., 2019) and Transformer methods (Gauthier & Levy, 2019; Toneva & Wehbe, 2019; Schwartz et al., 2019; Schrimpf et al., 2021; Antonello et al., 2021; Oota et al., 2022b; Aw & Toneva, 2023). Popular word embedding methods include textual (i.e., Word2Vec (Mikolov et al., 2013), fastText (Bojanowski et al., 2017), and GloVe (Pennington et al., 2014)), linguistic (i.e., dependency), conceptual (i.e., RWSGwn (Goikoetxea et al., 2015) and ConceptNet (Speer et al., 2017)), contextual (i.e., ELMo (Peters et al., 2018)). Popular sentence embedding models include average, max, concat of avg and max, SIF (Arora et al., 2017), SkipThoughts (Kiros et al., 2015), GenSen (Subramanian et al., 2018), InferSent (Conneau et al., 2017), ELMo, BERT (Devlin et al., 2019), RoBERTa (Liu et al., 2019), USE (Cer et al., 2018), QuickThoughts (Logeswaran & Lee, 2018) and GPT-2 (Radford et al., 2019). Transformer-based methods include pretrained BERT with various NLU tasks, finetuned BERT, Transformer-XL (Dai et al., 2019), GPT-2, BART (Lewis et al., 2020), BigBird (Zaheer et al., 2020), Longformer (Beltagy et al., 2020), and LongT5 (Guo et al., 2022). Experiential attribute models represent words in terms of human ratings of their degree of association with different attributes of experience, typically on a scale of 0-6 (Anderson et al., 2019; 2020; Berezutskaya et al., 2020; Just et al., 2010; Anderson et al., 2017b) or binary (Handjaras et al., 2016; Wang et al., 2017).
Figure 4: Context representation of several word orders: Past/Future context is constructed by considering words preceding/succeeding the current word (see Past/Future context illustrated for the current word vehicle for various orders).
Figure 5: Extraction of image representations: Prior research has explored the impact of various layer-wise image representations from CNN models (Yamins et al., 2014; Horikawa & Kamitani, 2017), for both brain encoding and decoding models. The plot of the image feature extraction was derived from the study by (Horikawa & Kamitani, 2017).
In the practice of employing word embeddings, encoding studies often utilize the average word representations within a given context or derive complete sentence representations through sentence embedding models. More recently, brain encoding research has shifted towards the use of contextualized word representations, examining how the amount of context affects the brain predictivity (Jain & Huth, 2018; Toneva & Wehbe, 2019). To obtain these contextualized word representations, Figure 4 illustrates how Past/Future context is constructed by considering words preceding/succeeding the current word. Given the constrained context length, each word is successively input to the network with at most $C$ previous tokens. For instance, given a story of $M$ words and considering the context length of 20, while the third word’s vector is computed by inputting the network with (w1, w2, w3), the last word’s vectors wM is computed by inputting the network with (wM-20, $\dots$, wM).
@@
Figure 6: Extraction of contextualized speech representations: Representation of the last frame within each window allows for the capture of temporal dynamics and contextual nuances in the speech signal. The length of the time window is typically varied from 16, 32, to 64 secs, with strides ranging from 10 to 100 milliseconds.
Visual stimulus representations. For visual stimuli, older methods used visual field filter bank (Thirion et al., 2006; Nishimoto et al., 2011) and Gabor wavelet pyramid (Kay et al., 2008; Naselaris et al., 2009). As shown in Figure 5, recent methods use models like CNNs (Du et al., 2020; Beliy et al., 2019; Anderson et al., 2017a; Yamins et al., 2014; Nishida et al., 2020) and concept recognition models (Anderson et al., 2020).
Audio stimuli representations. For audio stimuli, phoneme rate and presence of phonemes have been leveraged (Huth et al., 2016). Further, low-level speech features like filter banks (FBank), Mel Spectrogram, and MFCC from raw audio files, phonological features, articulation, and power spectrum (PowSpec) feature vectors were used in (Deniz et al., 2019). Recently, Nishida et al. (2020) used features from an audio deep learning model called SoundNet for audio stimuli representation. To extract representations from Transformer-based speech models such as Wav2Vec2.0, HuBERT and Whisper, Vaidya et al. (2022); Antonello et al. (2024); Oota et al. (2024a) varied the length of the time windows from 16, 32, to 64 seconds, with strides ranging from 10 to 100 milliseconds, as illustrated in Figure 6. Moreover, these studies utilized an autoregressive approach to derive speech representations. This method involves considering the representations of the last frame within each window, allowing for the capture of temporal dynamics and contextual nuances in speech.
Multimodal stimulus representations. To jointly model the information from multimodal stimuli, recently, various multimodal representations have been used. These include processing videos using audio+image representations like VGG (Simonyan & Zisserman, 2015) and SoundNet (Aytar et al., 2016) in (Nishida et al., 2020) or using image+text combination models like GloVe+VGG and ELMo+VGG in (Wang et al., 2020). Recently, the usage of multimodal text+vision models like Contrastive Language-Image Pretraining (CLIP) (Radford et al., 2021), Learning Cross-Modality Encoder Representations from Transformers (LXMERT) (Tan & Bansal, 2019), and VisualBERT (Li et al., 2020) was proposed in (Oota et al., 2022e).
@@ 3 Naturalistic Neuroscience Datasets
In this section, we discuss the popular text, visual, audio, video, and other multimodal neuroscience datasets that have been proposed in the literature. Table 1 shows a detailed overview of brain recording type, language, stimulus, number of subjects (||||S||||), and the task across datasets of different modalities. Figure 3 shows examples from a few datasets. A sample of naturalistic datasets is available at this link  [^1] https://neuroscout.org/datasets.
@@
Text Harry Potter (Wehbe et al., 2014) fMRI/ MEG English Reading Chapter 9 of Harry Potter and the Sorcerer’s Stone 9 Story understanding
240 Sentences with Content Words (Anderson et al., 2019) fMRI English Reading 240 active voice sentences describing everyday situations 14 Passive reading
Visual - (Thirion et al., 2006) fMRI - Viewing rotating wedges (8 times), expanding/contracting rings (8 times), rotating 36 Gabor filters (4 times), grid (36 times) 9 Passive viewing
Algonauts  (Cichy et al., 2019) fMRI/ MEG - Viewing 92 silhouette object images and 118 images of objects on natural background 15 Passive viewing
THINGS  (Hebart et al., 2023) fMRI/ MEG - Viewing 31188 natural images 8 Oddball Detection
@@
- (Brennan & Hale, 2019) EEG English Listening Chapter one of Alice’s Adventures in Wonderland (2,129 words in 84 sentences) as read by Kristen McQuillan 33 Question answering
@@
Video BBC’s Doctor Who  (Seeliger et al., 2019) fMRI English Viewing spatiotemporal visual and auditory videos (30 episodes). 120.8 whole-brain volumes ($\sim$23 h) of single-presentation data, and 1.2 volumes (11 min) of repeated narrative short episodes. 22 repetitions 1 Passive viewing
@@
Pippi Langkous  (Berezutskaya et al., 2020) ECoG Swedish/ Dutch Viewing 30 s excerpts of a feature film (in total, 6.5 min long), edited together for a coherent story 37 Passive viewing
@@
NATVIEW_EEGFMRI  (Telesford et al., 2023) fMRI/ EEG English Watching 5 short-length movies 22 Passive viewing
Other Multimodal 60 Concrete Nouns (Mitchell et al., 2008) fMRI English Viewing 60 different word-picture pairs from 12 categories, 6 times each 9 Passive viewing
@@
Pereira (Pereira et al., 2018) fMRI English Viewing 180 Words with Picture, Sentences, word clouds; reading 96 text passages; 72 passages. 3 times repeated. 16 Passive viewing and reading
@@
Table 1: Naturalistic Neuroscience Datasets. Publicly available datasets are linked to their sources in the Dataset column. In this table, |S| represents the number of participants in each dataset.
Text datasets. These datasets are created by presenting words, sentences, passages, or chapters as stimuli. Some of the text datasets include Harry Potter Story (Wehbe et al., 2014), ZuCo EEG (Hollenstein et al., 2018) and datasets proposed in (Handjaras et al., 2016; Anderson et al., 2017a; 2019; Wehbe et al., 2014). In Handjaras et al. (2016), participants were asked to verbally enumerate in one minute the properties (features) that describe the entities the words refer to. There were four groups of participants: 5 sighted individuals were presented with a pictorial form of the nouns, 5 sighted individuals with a verbal-visual (i.e., written Italian words) form, 5 sighted individuals with a verbal auditory (i.e., spoken Italian words) form, and 5 congenitally blind with a verbal auditory form. Data proposed by Anderson et al. (2017a) contains 70 Italian words taken from seven taxonomic categories (abstract, attribute, communication, event/action, person/social role, location, object/tool) in the law and music domain. The word list contains concrete as well as abstract words. ZuCo dataset (Hollenstein et al., 2018) contains sentences for which EEG recordings were obtained for 3 tasks: normal reading of movie reviews, normal reading of Wikipedia sentences, and task-specific reading of Wikipedia sentences. For this dataset curation, sentences were presented to the subjects in a naturalistic reading scenario. A complete sentence is presented on the screen. Subjects read each sentence at their own speed, i.e., the reader determines for how long each word is fixated and which word to fixate next.
Visual datasets. Older visual datasets were based on binary visual patterns (Thirion et al., 2006). Recent datasets contain natural images. Examples include Vim-1 (Kay et al., 2008), BOLD5000 (Chang et al., 2019), Algonauts (Cichy et al., 2019), NSD (Allen et al., 2022), Things-data (Hebart et al., 2023), NOD (Gong et al., 2023), and the dataset proposed in (Horikawa & Kamitani, 2017). BOLD5000 includes $\sim$20 hours of MRI scans per each of the four participants. 4,916 unique images were used as stimuli from 3 image sources. Algonauts contains two sets of training data, each consisting of an image set and brain activity in RDM format (for fMRI and MEG). Training set 1 has 92 silhouette object images, and training set 2 has 118 object images with natural backgrounds. Testing data consists of 78 images of objects on natural backgrounds. Most of the visual datasets involve passive viewing, but the dataset in (Horikawa & Kamitani, 2017) involved the participant doing the one-back repetition detection task.
@@
Audio datasets. Most of the proposed audio datasets are in English (Huth et al., 2016; Brennan & Hale, 2019; Anderson et al., 2020; Nastase et al., 2020), while there is one (Handjaras et al., 2016) on Italian, and another one (Li et al., 2021) in Chinese and French. The participants were involved in a variety of tasks while their brain activations were measured: Property generation (Handjaras et al., 2016), passive listening (Huth et al., 2016; Nastase et al., 2020), question answering (Brennan & Hale, 2019) and imagining themselves personally experiencing common scenarios (Anderson et al., 2020). In the last one, participants underwent fMRI as they reimagined the scenarios (e.g., resting, reading, writing, bathing, etc.) when prompted by standardized cues. Narratives (Nastase et al., 2020) used 27 different stories as stimuli. Across subjects, it is 6.4 days worth of recordings.

### v3

@@ 2 Stimulus Representations
This section discusses various stimulus representations proposed in the literature across different modalities: text, visual, audio, video, and multimodal stimuli.
Text stimulus representations. The evolution of text-based stimulus representations reflects the rapid advancements in natural language processing. Early methods relied on text corpus co-occurrence counts (Mitchell et al., 2008; Pereira et al., 2013; Huth et al., 2016), topic models (Pereira et al., 2013), and syntactic and discourse features (Wehbe et al., 2014). In recent years, both semantic models and experiential attribute models have gained prominence. Semantic representation models encompass a wide range of techniques, from word embeddings (Pereira et al., 2018; Wang et al., 2020; Pereira et al., 2016; Toneva & Wehbe, 2019; Anderson et al., 2017a; Oota et al., 2018) to more sophisticated sentence representation models (Sun et al., 2020; 2019; Toneva & Wehbe, 2019). The field has seen a shift from static word embeddings to contextualized representations, with models like RNNs (Jain & Huth, 2018; Oota et al., 2019) and Transformer-based methods (Gauthier & Levy, 2019; Toneva & Wehbe, 2019; Schwartz et al., 2019; Schrimpf et al., 2021; Antonello et al., 2021; Oota et al., 2022b; Aw & Toneva, 2023) gaining traction. Experiential attribute models offer a different perspective, representing words based on human ratings of their associations with various experiential attributes (Anderson et al., 2019; 2020; Berezutskaya et al., 2020; Just et al., 2010; Anderson et al., 2017b). These models provide insights into how humans conceptualize and experience language.
Figure 4: (a) Context Representation of Words in Language Models. This figure illustrates how past and future context is constructed for different word orders. Using the word “vehicle” as an example, we demonstrate how preceding words (past context) and succeeding words (future context) are considered for various context lengths. (b) Extraction of Image Representations for Brain-Computer Interface Models. This figure illustrates the process of extracting layer-wise image representations from Convolutional Neural Network (CNN) models. These representations have been extensively studied in prior research (Yamins et al., 2014; Horikawa & Kamitani, 2017) for their effectiveness in both brain encoding (predicting neural responses from visual stimuli) and decoding (reconstructing visual stimuli from brain activity) models. The figure demonstrates how different layers of a CNN, from early layers capturing low-level features to deeper layers representing more abstract concepts, can be utilized to understand and predict brain responses to visual stimuli. Figure is adapted from (Horikawa & Kamitani, 2017), and used with permission from the respective authors.
Recent brain encoding research has increasingly focused on contextualized word representations, examining how the amount of context affects brain predictivity (Jain & Huth, 2018; Toneva & Wehbe, 2019). Figure 4 (a) illustrates the construction of past/future context for these representations, demonstrating how researchers consider words preceding and succeeding the current word to create rich, context-aware embeddings.
In practice, the application of contextualized word representations often involves a constrained context length. This approach balances computational efficiency with the need for rich contextual information. For example, when processing a narrative of $M$ words with a context window of 20, the representation for each word is computed using up to 20 preceding words. Specifically, the vector for the third word would be derived from the input sequence $(w_{1},w_{2},w_{3})$, while the vector for the final word, $w_{M}$, would be based on the input $(w_{M-20},\ldots,w_{M})$. This sliding window approach allows the model to capture local context effectively while maintaining a manageable computational load, even for longer texts. Such methods enable researchers to investigate how varying amounts of context influence the model’s ability to predict brain activity, providing insights into the temporal aspects of language processing in the brain.
@@
Figure 5: Extraction of contextualized speech representations: Representation of the last frame within each window allows for the capture of temporal dynamics and contextual nuances in the speech signal. The length of the time window is typically varied from 16 to 64 secs, with strides ranging from 10 to 100 milliseconds.
Visual stimulus representations. The field of visual stimulus representation has similarly progressed from simpler methods to more complex, neural network-based approaches. Early techniques employed visual field filter banks (Thirion et al., 2006; Nishimoto et al., 2011) and Gabor wavelet pyramids (Kay et al., 2008; Naselaris et al., 2009). These methods provided a foundation for understanding how the brain processes visual information.
More recent approaches leverage the power of Convolutional Neural Networks (CNNs) (Du et al., 2020; Beliy et al., 2019; Anderson et al., 2017a; Yamins et al., 2014; Nishida et al., 2020) and concept recognition models (Anderson et al., 2020). CNNs, in particular, have revolutionized visual stimulus representation by automatically learning hierarchical features from raw image data. Figure 4 (b) illustrates how these CNN models are utilized for visual stimulus representation, showcasing their ability to capture increasingly abstract visual concepts across layers.
Audio stimuli representations. Audio stimulus representations have evolved from basic acoustic features to sophisticated deep learning models. Initial approaches focused on phoneme-level features (Huth et al., 2016) and low-level speech characteristics such as filter banks, Mel spectrograms, and MFCCs (Deniz et al., 2019). These methods provided a solid foundation for understanding speech processing in the brain.
Recent advancements have seen the integration of deep learning models specifically designed for audio processing. For instance, Nishida et al. (2020) employed features from SoundNet, a deep neural network trained on large-scale audiovisual data. The field has further progressed with the adoption of Transformer-based speech models like Wav2Vec2.0, HuBERT, and Whisper (Vaidya et al., 2022; Antonello et al., 2024b; Oota et al., 2024a). These models offer rich, contextual representations of speech, capturing nuances that were previously difficult to model. Figure 5 demonstrates how researchers extract representations from these models using various time window lengths and strides, allowing for a more comprehensive analysis of temporal dynamics in speech processing.
Multimodal stimulus representations. The increasing focus on naturalistic stimuli has led to significant developments in multimodal representations. These approaches aim to jointly model information from multiple sensory inputs, mirroring the integrative nature of human perception. Recent studies have explored various combinations of modalities, each offering unique insights into how the brain processes complex, real-world stimuli.
For video processing, researchers have combined audio and image representations, such as using VGG for visual features alongside SoundNet for audio features (Nishida et al., 2020). This approach allows for a more comprehensive understanding of how the brain integrates audiovisual information. In the realm of image and text combinations, models like GloVe+VGG and ELMo+VGG (Wang et al., 2020) have been employed, leveraging the strengths of both visual and linguistic representations.
The latest advancements in multimodal representations include sophisticated models that are pretrained on large-scale multimodal datasets. These include Contrastive Language-Image Pretraining (CLIP) (Radford et al., 2021), Learning Cross-Modality Encoder Representations from Transformers (LXMERT) (Tan & Bansal, 2019), and VisualBERT (Li et al., 2020; Oota et al., 2022e). These models offer powerful, joint representations of visual and linguistic information, potentially providing deeper insights into how the brain integrates information across modalities.
@@ 3 Naturalistic Neuroscience Datasets
In this section, we discuss popular neuroscience datasets that involve audio, visual, and other multimodal stimuli that have been utilized in the literature. Tables 2, 3 and 4 provide detailed overview of types of brain recording, languages used, stimuli presented, number of subjects (||||S||||), and tasks across datasets of different modalities. The audio datasets consist of listening to narrative stories or spoken utterances, while the vision-based datasets are categorized into reading stories and watching static images or silent videos. The multimodal stimulus datasets include the presence of visual input along with other modalities, such as audio. Figure 3 illustrates examples from selected datasets.
Table 2: Naturalistic Neuroscience Audio Datasets. Publicly available datasets are linked to their sources in the Dataset column. In this table, |S| represents the number of participants in each dataset.
@@
Blank2014 (Blank et al., 2014) fMRI English Listening naturalistic stories 10 Passive listening
@@
Alice (Brennan & Hale, 2019) EEG English Listening Chapter one of Alice’s Adventures in Wonderland (2,129 words in 84 sentences) as read by Kristen McQuillan 33 Question answering
@@
SMN4Lang  (Wang et al., 2022) fMRI, MEG Chinese Listening 6 hours of naturalistic stories 12 Passive listening
Martin2023  (Martin, 2023) MEG Dutch, French Listening 70 mins of naturalistic stories 25 Passive listening
Giovanni2023  (Liberto et al., 2023) EEG English Listening an audio book: “The Old Man and the Sea” 19 Passive listening
Little Prince Hong Kong  (Momenian et al., 2024) fMRI, EEG Cantonese Listening audiobook for about 20 minutes. 52 Passive listening
SCNU-Mandarin-Cantonese  (Yuan & Yang, 2024) fMRI Mandarin, Cantonese Listening three narratives 27 Passive listening
Table 3: Naturalistic Neuroscience Vision-based Datasets (Reading text, Watching images and videos (without audio). Publicly available datasets are linked to their sources in the Dataset column. In this table, |S| represents the number of participants in each dataset.
Dataset Authors Type Lang. Stimulus ||||S|||| Task
Text Harry Potter (Wehbe et al., 2014) fMRI, MEG English Reading Chapter 9 of Harry Potter and the Sorcerer’s Stone 9 Story understanding
Fedorenko2016 (Fedorenko et al., 2016) ECoG English 8-word-long sentences 5 Passive reading
Pereira (Pereira et al., 2018) fMRI English Viewing 180 Words with Picture, Sentences, word clouds; reading 96 text passages; 72 passages. 3 times repeated. 16 Passive reading
- (Anderson et al., 2019) fMRI English Reading 240 active voice sentences describing everyday situations 14 Passive reading
Greta2024  (Tuckute et al., 2024) fMRI English Reading 1,000 sentences 9 Passive reading
Static Images Inverse retinotopy (Thirion et al., 2006) fMRI - Viewing rotating wedges (8 times), expanding/contracting rings (8 times), rotating 36 Gabor filters (4 times), grid (36 times) 9 Passive viewing
Algonauts  (Cichy et al., 2019) fMRI, MEG - Viewing 92 silhouette object images and 118 images of objects on natural background 15 Passive viewing
THINGS  (Hebart et al., 2023) fMRI, MEG - Viewing 31188 natural images 8 Oddball Detection
Gifford2022 EEG  (Gifford et al., 2022) EEG - Viewing 82160 natural images 8 Passive viewing
Infants EEG  (Quek et al., 2024) EEG - Viewing 200 natural objects 42 Passive viewing
Videos (without audio) BBC’s Doctor Who  (Seeliger et al., 2019) fMRI English Viewing spatiotemporal visual and auditory videos (30 episodes). 120.8 whole-brain volumes ($\sim$23 h) of single-presentation data, and 1.2 volumes (11 min) of repeated narrative short episodes. 22 repetitions 1 Passive viewing
@@
Pippi Langkous  (Berezutskaya et al., 2020) ECoG Swedish, Dutch Viewing 30 s excerpts of a feature film (in total, 6.5 min long), edited together for a coherent story 37 Passive viewing
@@
NATVIEW_EEGFMRI  (Telesford et al., 2023) fMRI, EEG English Watching 5 short-length movies 22 Passive viewing
Mind captioning  (Horikawa, 2024) fMRI English Watching total of 2,196 videos 6 Passive viewing
Table 4: Naturalistic Neuroscience Multimodal Datasets. Publicly available datasets are linked to their sources in the Dataset column. In this table, |S| represents the number of participants in each dataset.
Dataset Authors Type Lang. Stimulus ||||S|||| Task
Multimodal 60 Concrete Nouns (Mitchell et al., 2008) fMRI English Viewing 60 different word-picture pairs from 12 categories, 6 times each 9 Passive viewing
@@
fMRI Narrative movie  (Yamaguchi et al., 2024) fMRI English Watching nine narrative movies 6 Passive viewing and listening
Game of Thrones  (Watson & Andrews, 2024) fMRI English Watching short audiovisual clips 73 Passive viewing and listening
AVID  (Sava-Segal et al., 2023) fMRI English Watching short audiovisual clips 48 Passive viewing and listening
3.1 Audio datasets
@@
Most proposed audio datasets are in English (Huth et al., 2016; Brennan & Hale, 2019; Anderson et al., 2020; Nastase et al., 2020), with one in Italian (Handjaras et al., 2016) and another in Chinese and French (Li et al., 2021). Participants engaged in various tasks while their brain activations were measured, including property generation (Handjaras et al., 2016), passive listening (Huth et al., 2016; Nastase et al., 2020), question answering (Brennan & Hale, 2019), and imagining themselves personally experiencing common scenarios (Anderson et al., 2020). In the latter, participants underwent fMRI as they reimagined scenarios (e.g., resting, reading, writing, bathing) when prompted by standardized cues. The Narratives dataset (Nastase et al., 2020) used 27 different stories as stimuli, totaling 6.4 days worth of recordings across subjects.
3.2 Vision datasets
Text datasets: Reading. These datasets are created by presenting words, sentences, passages, or chapters as stimuli. Examples include the Harry Potter Story dataset (Wehbe et al., 2014), ZuCo EEG dataset (Hollenstein et al., 2018), and datasets proposed in (Handjaras et al., 2016; Anderson et al., 2017a; 2019; Wehbe et al., 2014). In (Handjaras et al., 2016), participants were asked to verbally enumerate, within one minute, the properties (features) describing the entities the words refer to. The study involved four groups of participants: 5 sighted individuals presented with pictorial forms of nouns, 5 sighted individuals with written Italian words, 5 sighted individuals with spoken Italian words, and 5 congenitally blind individuals with spoken Italian words. The dataset proposed by Anderson et al. (2017a) contains 70 Italian words from seven taxonomic categories (abstract, attribute, communication, event/action, person/social role, location, object/tool) in the law and music domains, including both concrete and abstract words. The Pereira dataset includes pictures, sentences, and word clouds, for which fMRI recordings were obtained during three viewing conditions (Pereira et al., 2018). The ZuCo dataset (Hollenstein et al., 2018) contains sentences for which EEG recordings were obtained during three tasks: normal reading of movie reviews, normal reading of Wikipedia sentences, and task-specific reading of Wikipedia sentences. For this dataset, sentences were presented to subjects in a naturalistic reading scenario, with each complete sentence displayed on the screen. Subjects read each sentence at their own pace, determining how long to fixate on each word and which word to fixate on next.
Static Image datasets. Earlier visual datasets were based on binary visual patterns (Thirion et al., 2006). Recent datasets contain natural images, including Vim-1 (Kay et al., 2008), BOLD5000 (Chang et al., 2019), Algonauts (Cichy et al., 2019), NSD (Allen et al., 2022), Things-data (Hebart et al., 2023), NOD (Gong et al., 2023), and the dataset proposed in (Horikawa & Kamitani, 2017). BOLD5000 includes approximately 20 hours of MRI scans per participant for four participants, using 4,916 unique images as stimuli from three image sources. The Algonauts dataset contains two sets of training data, each consisting of an image set and brain activity in RDM format (for fMRI and MEG). Training set 1 has 92 silhouette object images, while training set 2 has 118 object images with natural backgrounds. The testing data consists of 78 images of objects on natural backgrounds. Most visual datasets involve passive viewing, but the dataset in (Horikawa & Kamitani, 2017) involved participants performing a one-back repetition detection task.